OpenStudy (anonymous):

Add. 12.6 + (–12.6) A. 25.2 B. 24.2 C. 0 D. –25.2

OpenStudy (alexandervonhumboldt2):

12.6+(-12.6)=12.6-12.6=simplify

OpenStudy (alexandervonhumboldt2):

x+(-a)=x-a

OpenStudy (anonymous):

0 is the answer